Β (a)How do you shield a given region from any external electrostatic field? Explain briefly.

(b) Determine a unit vector perpendicular to the plane of the two forces 𝐴⃗= πš€Μ‚+2 πš₯Μ‚+ 3π‘˜^

and B-> =5πš€Μ‚+10πš₯Μ‚+ 2π‘˜^


1
Expert's answer
2021-05-20T10:11:38-0400

a) Electrostatic shielding is defined as a phenomenon that is seen when a Faraday cage is used to block the effects of an electric field. The effects of external fields on the internal contents are blocked using the cage.

b)

"\vec{A}\times\vec{B}=(1,2,3)\times(5,10,2)=(-26,13,0)\\\vec{e}=\frac{1}{\sqrt{5}}(-2,1,0)"